# MonoSOWA: Scalable monocular 3D Object detector Without human Annotations

### v1

@@ Abstract
Detecting the three-dimensional position and orientation of objects using a single RGB camera is a foundational task in computer vision with many important applications. Traditionally, 3D object detection methods are trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with the ever-increasing amounts of data being captured.
In this paper, we present the first method to train 3D object detectors for monocular RGB cameras without domain-specific human annotations, thus making orders of magnitude more data available for training. Thanks to newly proposed Canonical Object Space, the method can not only exploit data across a variety of datasets and camera setups to train a single 3D detector, but unlike previous work it also works out of the box in previously unseen camera setups. All this is crucial for practical applications, where the data and cameras are extremely heterogeneous.
The method is evaluated on two standard autonomous driving datasets, where it outperforms previous works, which, unlike our method, still rely on 2D human annotations. The source code and the model is going to be published soon.
@@ I Introduction
Monocular 3D object detection is a key component of many important systems, ranging from robotics to autonomous cars. 3D object detection methods are typically trained in a fully-supervised setup, requiring vast amounts of human annotations, which are laborious, costly, and do not scale well with ever-increasing amounts of data that ideally should be used for training. To make matters even worse, especially in autonomous driving, new cameras and new camera setups emerge constantly, which typically implies the need to collect, label and train on a new dataset as camera parameters such as focal length change.
In this paper, we address both of these challenges – costly human labels and changing camera setup – by introducing a new method to train a 3D rigid object detector in a Canonical Object Space without domain-specific human labels. Our method allows us to exploit large amounts of training data readily available because all we need as input for training are unlabelled driving video sequences. Furthermore, our method does not require data from additional sensors such as LiDAR. It makes our method extremely applicable in practice because as such our method can directly exploit data collected by a majority of currently manufactured vehicles, which typically only have a single camera. This is in contrast with some methods which require LiDAR (and human annotations) [1, 2], because for them specialised vehicles are needed to collect the data, and this severely limits the variety of geographies, environments, and traffic situations being captured.
Drawing inspiration from monocular depth literature [3], the proposed Canonical Object Space (COS) allows us to aggregate and learn from data coming from different cameras, which is a crucial pre-condition to being able to effectively harvest otherwise very heterogeneous video sequences. Moreover, unlike previous methods such as VSRD [4], it also allows for the deployment of our 3D object detector in new unseen camera setup without needing to train or fine-tune for the new setup. This is again extremely useful in practice, because new camera sensors and setups are developed all the time, but it is unfeasible to collect data, label them by human annotators, and then train a new model for each new camera type.
@@
    We propose a novel auto-labelling pipeline which exploits temporal consistency in video sequences to automatically generate 3D position, size, and orientation labels of rigid objects (cars) from a monocular camera.
@@
    To the best of our knowledge, our method is the first method to train a monocular 3D object detector for autonomous driving which does not require neither 2D nor 3D human annotations; we demonstrate to the community that such an approach is feasible and is worth pursuing further.
@@
    We propose novel Canonical Object Space (COS), which allows us to a) aggregate data from different camera setups in training, and b) at inference time, use an already trained model in the previously unseen camera setup, i.e. without any training or fine-tuning for the new setup.
  4. 4.
    Our auto-labelling method is significantly faster than previous work, allowing us to process significantly more training data than the current state-of-the-art method.
@@ II Related work
Monocular Depth Estimation
@@
Monocular depth estimation has shown significant improvements in recent years. MiDAS [5] employs affine-invariant losses to enable training on multiple datasets. ZoeDepth [6] combines training on multiple datasets with relative losses and further fine-tunes on the target metric dataset. Depth Anything [7] focuses on the affine-invariant depth estimation using self-supervised pre-training. For metric depth estimation, which is relevant for object detection, it fine-tunes on the target dataset such as KITTI [8]. Metric3D [9, 3] focuses on metric depth estimation as it uses a Canonical Camera Transformation Module to further improve zero-shot performance, together with DINOv2 [10] or ConvNext [11] as the backbone and DPT [12] as the predictor. CamConv [13] models cameras inside the network, but as shown in [3], this leads to poor results.
Fully-supervised Monocular 3D Object Detection
@@
Early methods [14, 15, 16] were based on lifting the image into pseudo-LiDAR and then detecting objects in the generated point cloud. SMOKE [17], on the other hand, directly predicts 3D bounding boxes from images. MonoDTR [18] and MonoDETR [19] use an end-to-end depth-aware/guided transformer [20] architecture to directly predict 3D bounding boxes. SSD-MonoDETR [21] introduces scale awareness into the model. MonoATT [22] variably assigns tokens to areas of higher importance. MonoFlex [23] uses an ensemble of direct depth and keypoint-based predictions and decouples the detection of truncated and untruncated vehicles. MonoCD [24] enhances the ensemble with a complementary depth predictor. However, all the aforementioned methods require 3D human annotations during their training.
Weakly-supervised 3D Object Detection
Most previous works on weakly-supervised 3D detection have focused on training 3D detectors for LiDAR. VS3D [25] pioneered in this area by using point density in LiDAR scans to identify objects, and detections are further refined by a student-teacher network, where the teacher is trained on images only. Zakharov et al. [2] employs a 2D off-the-shelf detector accompanied by a novel Signed Distance Fields (SDF)-renderer within the DeepSDF framework [26] and Normalized Object Coordinate Space (NOCS). The method is trained on a synthetic dataset and further fine-tuned on real datasets.
@@
WS3D [27, 28] employs Bird’s Eye View (BEV) human-centered clicks as weak supervision to train a network for detecting vehicles, arguing that these are much easier to obtain than 3D bounding boxes. Furthermore, they use a small subset of human ground-truth annotations to train a second network to predict the spatial dimensions of the detected vehicles. McCraith et al. [29] employ direct fitting of generic templates into LiDAR scans segmented by Mask-RCNN [30] and share the information between frames by requiring consistency between frames while simultaneously discounting outliers. TCC-det [31] uses direct fitting of generic templates on aggregated points corresponding to each instance. During the training of the 3D detector, TCC-det appends two additional losses for finer alignment. VG-W3D [32] encodes the image and LiDAR scan separately and then enforces alignment between 2D and 3D on multiple feature levels to predict precise 3D bounding boxes.
@@
Figure 2: Our auto-labelling pipeline creates 3D pseudo-groundtruth using only images, camera parameters and IMU data as inputs.
@@
In contrast to the aforementioned methods, VSRD [4] focuses on 3D detection in camera, and does not require LiDAR scans. However, it uses 2D human ground-truth masks for each instance, representing them as a surface in Signed Distance Fields (SDF). Furthermore, the SDF are rendered into masks through their proposed volumetric-aware silhouette rendering, enabling end-to-end training. Each instance is optimized over multiple frames. While VSRD shows promising results in both weakly-supervised and semi-supervised training, it requires 2D instance masks, and rendering of each frame takes approximately 15 minutes. In contrast, our method does not require any human annotations at all and is approximately 100 times faster.
@@ III Method
Traditionally, deep networks that predict 3D bounding boxes of vehicles (and other agents) from a single image are trained using 3D annotations, which are manually created by human annotators. They typically rely on both camera image and LiDAR data for correct distance estimate in creating their labels. However, this process is very laborious and costly, limiting the amount of training data available to train such methods. Moreover, it requires capturing the scene using both a camera and a LiDAR sensor, which is a significant limitation since the majority of production vehicles these days do not have LiDAR on board – as a result, data collection is only limited to vehicles dedicated to such a purpose, and this again limits the variety and diversity of data being used to train these methods.
Our method overcomes these limitations by introducing a novel approach to train a 3D object detector using only images (and vehicle motion data) as inputs, without requiring human labels or LiDAR in the training process.
III-A Auto-labelling pipeline
The process of automatic label creation assumes that images are captured in a sequence, which is a natural way to capture data in driving scenarios. It also assumes that the camera setup is known (both intrinsics and extrinsics) for each sequence. However, the camera setup might be and actually often is different (see section III-B). Lastly, it assumes approximate data about ego-vehicle’s motion is available (e.g. from GPS or an Inertial Measurement Unit – IMU), which again is commonplace.
Because labelling a 7-DOF pose of objects jointly is a very challenging problem, we instead infer individual pose parameters separately in a sequence of steps.
Detecting objects in pseudo-LiDAR
The auto-labelling process begins by inferring a metric depth map $D\in\mathbb{R}^{h\times w}$ for each camera image $I\in\mathbb{R}^{3\times h\times w}$, where $h$ and $w$ denotes image height and width, using an off-the-shelf monocular depth estimator. We opted to use Metric3D [3], as it is zero-shot, it does not limit our method in terms of data distribution and has shown great results for generalization. The choice of a particular monocular metric depth estimator is, however, not crucial. For each pixel and its corresponding depth, a 3D point cloud is generated as
$X_{u,v}$ $=\frac{D_{u,v}\cdot(u-c_{x})}{f_{x}}$
$Y_{u,v}$ $=\frac{D_{u,v}\cdot(v-c_{y})}{f_{y}}$
@@ III-A Auto-labelling pipeline
where $u\in[0,h]$ and $v\in[0,w]$, $c$ is the principal point and $f$ is thefocal length. We denote the pseudo lidar point cloud as $P_{i}\in\mathbb{R}^{3\times hw}$, where $i$ stands for the frame index.
Figure 3: Combining instance segmentation in image (top) with monocular metric depth estimation (middle) enables us to produce pseudo-LiDAR point cloud for each object.
Next, inspired by [29], we employ an off-the-shelf 2D object image detector to detect the objects we are interested in (vehicles) and their instance segmentation masks in each camera image. Given an instance segmentation mask $M_{i,j}\in\mathbb{R}^{h\times w}$ of an object $j$ in a frame $i$, we find object point cloud $P_{i,j}$ in the point cloud as
@@
where $\mathcal{T}(p)$ denotes the projection of the 3D point $p$ into image coordinates (using camera calibration, which is known).
The $P_{i,j}$ tends to contain outliers, as either the masks or the metric depth estimation is not perfect. To remove the outliers, an ensemble of the following methods is employed: HBOS [33], Z-scores, statistical outlier rejection in Open3D [34], HDBSCAN [35] and DBSCAN [36]. If at least two of these methods classify a point as an outlier, it is removed. To compensate for the effect of the ego-vehicle movement, we transform all $P_{i,j}$ into the reference frame coordinate space, given the poses from IMU data.
Temporal consistency
In order to exploit temporal consistency across multiple frames of the sequence, we need to first establish correspondences of individual object instances between frames. Because we have inferred approximate 3D point cloud of each object instance $P_{i,j}$ in each frame, we track each object in 3D world coordinate system.
For each object $j$ in frame $i$, we approximate its location $L_{i,j}$ as the median of its point cloud $P_{i,j}$. The tracking is initialized in $n$ frames before the reference (current) frame, and for each frame, it matches the instances based on locations in the current frame $L_{i,j}$ and the predicted future locations from the last frame into the current frame $\hat{L}_{i,j}$. For the location prediction $\hat{L}_{i,j}$, a simple physics-aware motion model is used
@@
To match two object instances between frames, it is required that the instances are both their nearest neighbours and that the distance between them is lower than a set threshold $T_{\text{dist}}$. Otherwise, the instance is considered as a new object. After this step, each object instance is represented by extracted points $P_{i,j}$ and locations $L_{i,j}$ in each frame the instance is present. The tracking sequentially processes all the frames up to $n$ after the reference frame.
Movement Classification
The next step is to classify instances as either stationary or moving, as different principles of temporal consistency exploitation are applied to those classes. Note that all object instances actually appear moving – they change location in relative terms, as the ego-vehicle itself is driving while data is captured.
In order to classify a vehicle as stationary or moving, we propose the following novel classification procedure. For each instance, all locations $L_{m...n,j}$ are taken and differences between all adjacent locations are computed, where $m$ is the first frame, where the instance is present and $n$ is the last. From those differences, the standard deviation $\sigma$ is calculated as
$\sigma_{j}$ $=\frac{1}{\sqrt{2}}\sqrt{\frac{1}{n-m}}\sum_{i=m+1}^{n}(\mu_{j}-D_{i,j})^{2}$ (4)
$\mu_{j}$ $=\frac{1}{n-m}\sum_{i=m+1}^{n}D_{i,j}\qquad D_{i,j}=L_{i,j}-L_{i-1,j}$ (5)
Furthermore, the actual distance $d$ and expected distance $\hat{d}$ caused by the $\sigma$ is calculated as
$d_{j}=\left\lVert L_{m,j}-L_{n,j}\right\rVert^{2}_{2}\qquad\hat{d}_{j}=(n-m)\cdot\left\lVert\sigma_{j}\right\rVert^{2}_{2}$ (6)
The $z$- and $p$-scores are calculated to know if the difference between the $d_{\text{real}}$ and $d_{\text{exp}}$ is statistically significant or not
$z_{j}=\frac{d_{j}}{\hat{d}_{j}}\qquad p_{j}=1-\Phi(z_{j})$ (7)
To classify the instance as moving, the $p$-score is required to be lower than the threshold $T_{m}$ and simultaneously, the $d$ distance must be higher than a given threshold.
As all $P_{i,j}$ are in the reference frame, we can aggregate points for stationary instances by simply concatenating the points
$A_{j}=\bigcup_{i=m}^{n}P_{i,j}$ (8)
Aggregating points for moving instances poses a significant challenge, as the $P_{i,j}$ differs significantly between frames. For that reason, for moving instances, only the trajectory and $P_{i,j}$ are employed.

### v2

@@ Abstract
Inferring object 3D position and orientation from a single RGB camera is a foundational task in computer vision with many important applications. Traditionally, 3D object detection methods are trained in a fully-supervised setup, requiring LiDAR and vast amounts of human annotations, which are laborious, costly, and do not scale well with the ever-increasing amounts of data being captured.
We present a novel method to train a 3D object detector from a single RGB camera without domain-specific human annotations, making orders of magnitude more data available for training. The method uses newly proposed Local Object Motion Model to disentangle object movement source between subsequent frames, is approximately 700 times faster than previous work and compensates camera focal length differences to aggregate multiple datasets.
The method is evaluated on three public datasets, where despite using no human labels, it outperforms prior work by a significant margin. It also shows its versatility as a pre-training tool for fully-supervised training and shows that combining pseudo-labels from multiple datasets can achieve comparable accuracy to using human labels from a single dataset. The source code and model will be published soon.
@@ I Introduction
Monocular 3D object detection is a key component of many important systems, ranging from robotics to autonomous cars. 3D object detection methods are typically trained in a fully-supervised setup, requiring specialized vehicles to simultaneously capture RGB image and LiDAR modalities, followed by human annotators manually creating 3D labels. This is extremely laborious, costly, and limits the diversity of training data being captured. To make matters even worse, especially in autonomous driving, new camera setups emerge constantly, which typically means this laborious and costly process needs to be repeated again.
In this paper, we address both of these challenges – costly human labelling and reliance on dedicated vehicles to capture data – by introducing a new method to train a monocular 3D rigid object detector without domain-specific human labels and without LiDAR. Our method allows us to exploit large amounts of training data readily available because it does not rely on human annotations, which traditionally is the main limiting factor in exploiting captured data, and it does not need additional sensors such as LiDAR, which implies our method can directly exploit data collected by a majority of currently manufactured vehicles, which typically only have a single camera. This is in contrast with some methods which require LiDAR (and human annotations) [1, 2] as specialized vehicles are needed to collect the data. This then severely limits the diversity of geographies, environments, and traffic situations captured.
In our method, we automatically create 3D object labels (7-DOF poses) by using outputs of two generic off-the-shelf models – a 2D object detector and a monocular depth estimator – and by combining them into a single pipeline that processes video sequences captured from a moving vehicle. The proposed pipeline uses temporal consistency of objects in real world to precisely infer object 3D object orientation, position and dimensions in a series of consecutive steps, using newly proposed Local Object Motion Model (LOMM). LOMM disentangles the position change between subsequent frames into ego-motion of the vehicle capturing the scene and motion of the labeled object itself, allowing our method, unlike previous work [3, 4], to aggregate information for both stationary and moving objects throughout the whole video sequence, which then results in more precise 3D orientation and position 3D label in every frame. To address the heterogeneity of captured data in practice, we are the first weakly-supervised method to address this issue by adopting Canonical Object Space (COS) [5, 6] to effectively harvest and auto-label video sequences captured by different camera setups.
@@
    To the best of our knowledge, our method is the first method to train a monocular 3D object detector for autonomous driving without 2D or 3D human annotations; we demonstrate to the community that such an approach is feasible and is worth pursuing further.
@@
    We propose a novel auto-labelling pipeline which exploits temporal consistency in video sequences to generate 3D position, size, and orientation labels of rigid objects (cars) from a monocular camera. The pipeline uses novel Local Object Motion Model (LOMM) to disentangle ego-motion of the vehicle capturing the scene and motion of the labelled object itself to aggregate information both for stationary and moving objects.
@@
    Our auto-labelling method is almost 700 times faster than previous work and is able to aggregate data captured by different camera setups, allowing us to process significantly more data than the current state-of-the-art methods [4], where their auto-labelling speed in practice prohibits applications to large-scale datasets.
@@ II Related work
Monocular Depth Estimation.
@@
Monocular depth estimation has shown significant improvements in recent years. MiDAS [7] employs affine-invariant losses to enable training on multiple datasets. ZoeDepth [8] combines training on multiple datasets with relative losses and further fine-tunes on the target metric dataset. Depth Anything [9] focuses on the affine-invariant depth estimation using self-supervised pre-training. For metric depth estimation, which is relevant for 3D object detection, it fine-tunes on the target dataset such as KITTI [10]. Metric3D [11, 5] focuses on metric depth estimation as it uses a Canonical Camera Transformation Module to further improve zero-shot performance, together with DINOv2 [12] or ConvNext [13] as the backbone and DPT [14] as the predictor. We opt to use Metric3D [11, 5] for its superior zero-shot performance.
Fully-supervised Monocular 3D Object Detection.
@@
Early methods [15, 16, 17] were based on lifting the image into pseudo-LiDAR and then detecting objects in the generated point cloud. SMOKE [18], on the other hand, directly predicts 3D bounding boxes from images. MonoDTR [19] and MonoDETR [20] use an end-to-end depth-aware/guided transformer [21] architecture to directly predict 3D bounding boxes. SSD-MonoDETR [22] introduces scale awareness into the model. MonoATT [23] variably assigns tokens to areas of higher importance. MonoFlex [24] uses an ensemble of direct depth and keypoint-based predictions and decouples the detection of truncated and untruncated vehicles. MonoCD [25] enhances the ensemble with a complementary depth predictor. However, all the aforementioned methods were previously trained using 3D human labels. We utilize MonoDETR [20] for its competitive performance and high-quality code implementation.
Weakly-supervised LiDAR 3D Object Detection.
@@
The majority of previous work on weakly-supervised 3D detection have focused on training 3D detectors for LiDAR. VS3D [26] pioneered in this area by using point density in LiDAR scans to identify objects. Zakharov et al. [2] employs a 2D off-the-shelf detector accompanied by a novel Signed Distance Fields (SDF)-renderer and Normalized Object Coordinate Space (NOCS). WS3D [27, 28] employs Bird’s Eye View (BEV) human-centred clicks as weak supervision to train a network for detecting vehicles. McCraith et al. [29] employ direct fitting of generic templates into LiDAR scans while requiring consistency between frames for stationary vehicles and discounting outliers. TCC-det [30] uses direct fitting of generic templates on aggregated information corresponding to each instance and trains with two additional losses for finer alignment. VG-W3D [31] encodes the image and LiDAR scan separately and then enforces alignment between 2D and 3D on multiple feature levels to predict precise 3D boxes.
@@
Figure 2: Our auto-labelling pipeline creates 3D labels using only images and ego-motion data as inputs (and known camera parameters).
Unsupervised LiDAR 3D Object Detection
Several approaches have been proposed for unsupervised object detection and tracking in LiDAR scans. Cheng et al. [32] focus on open-set 3D object detection using the MLUC network, which identifies objects while keeping some unclassified. ClusterNet [33] performs unsupervised object discovery using both 3D LiDAR and 2D RGB images. Najibi et al.  [34] involve scene-flow estimation and meta-label creation for detection and trajectory prediction. MODEST [35] detects objects via clustering and refines them in a self-training cycle. DRIFT [36] uses a Persistency prior score for noisy label creation and reinforced learning for fine-tuning. Recent approaches include density-based clustering and self-supervised learning for object detection [37], clustering based on the ”what moves together” paradigm [38], and self-supervised LiDAR flow estimation for motion cue generation [39]. While [32, 37] do not differentiate between moving and stationary vehicles, [33, 34, 35, 36, 38, 39] only detect moving vehicles and assume the detector will generalize to detect all movable vehicles (including stationary).
Weakly-supervised Monocular 3D Object Detection
@@
WeakMono3D [3] is the first method not using LiDAR scans, as only sequences of images with 2D ground-truth labels are required as input. VSRD [4] represents each instance as a surface in Signed Distance Fields (SDF). The SDFs are rendered into masks through their proposed volumetric-aware silhouette rendering, enabling end-to-end training. Each instance is optimized over multiple frames. Both WeakMono3D and VSRD struggle with moving vehicles, which are either discarded or given low confidence; on the other hand, our work can leverage information from both moving and stationary vehicles. While WeakMono3D and VSRD show promising results in both weakly-supervised and semi-supervised training, they require 2D instance masks, and in VSRD rendering of each frame takes approximately 15 minutes. In contrast, our method does not require any human annotations at all and is approximately 700 times faster.
@@ III Method
The process of automatic label creation assumes that images are captured in a sequence, which is a natural way to capture data in driving scenarios. It also assumes that the camera setup is known (both intrinsics and extrinsics) for each sequence, although each sequence might have different camera setup. Lastly, it assumes approximate data about ego-vehicle’s motion is available (e.g. from GPS or IMU), which again is commonplace. Because labelling a 7-DOF pose of objects jointly (position, size, yaw) is a very challenging problem, we instead infer individual pose parameters separately in a sequence of steps (see fig. 2).
III-A Pseudo-LiDAR aggregation
The auto-labelling process begins by inferring a metric depth map $D\in\mathbb{R}^{h\times w}$ for each camera image $I\in\mathbb{R}^{3\times h\times w}$, where $h$ and $w$ denotes image height and width, using an off-the-shelf monocular depth estimator. We opted to use Metric3D [5], as it is zero-shot, it does not limit our method in terms of data distribution and it has shown great generalization results.
Using the inferred depth map $D$, a 3D point cloud is generated as
$X_{u,v}$ $=\frac{D_{u,v}\cdot(u-c_{x})}{f_{x}}\qquad Y_{u,v}=\frac{D_{u,v}\cdot(v-c_{y})}{f_{y}}$
@@ III-A Pseudo-LiDAR aggregation
where $u\in[0,h]$ and $v\in[0,w]$, $c$ is the principal point and $f$ is the focal length. We denote the pseudo-LiDAR point cloud as $P_{i}\in\mathbb{R}^{3\times hw}$, where $i$ stands for the frame index.
Figure 3: Combining instance segmentation in image (top) with monocular metric depth estimation (middle) enables us to produce pseudo-LiDAR point cloud for each object (bottom).
Next, inspired by [29], we employ an off-the-shelf 2D object image detector [40] to detect the objects we are interested in (vehicles) and their instance segmentation masks in each camera image. Given an instance segmentation mask $M_{i,j}\in\mathbb{R}^{h\times w}$ of an object $j$ in a frame $i$, we find object point cloud $P_{i,j}$ in the point cloud as
@@
where $\mathcal{T}(p)$ denotes the projection of the 3D point $p$ into image coordinates, as shown in fig. 3.
Tracking.
In order to exploit temporal consistency of given object across the sequence frames, we need first to establish correspondences of the object instance between frames. Because we have inferred the approximate 3D point cloud of each object instance $P_{i,j}$ in each frame, we track each object in the 3D world coordinate system.
We approximate spatial location $L_{i,j}\in\mathbb{R}^{3}$ of each object $j$ in frame $i$ as the median of its point cloud $P_{i,j}$. The tracking is initialized in $n$ frames before the reference (current) frame, and for each frame, it matches the instances based on locations in the current frame $L_{i,j}$ and the predicted future locations from the last frame into the current frame $\hat{L}_{i,j}$. For the location prediction $\hat{L}_{i,j}$, a simple physics-aware motion model is used
@@
To match two object instances between frames, it is required that the instances are both their nearest neighbours and that the distance between them is lower than a set threshold $T_{\text{dist}}$. Otherwise, the instance is considered as a new object. After this step, each object instance is represented by extracted points $P_{i,j}$ and locations $L_{i,j}$ in each frame the instance is present. The tracking sequentially processes all the frames up to $m$ after the reference frame. Finally, we transform all $P_{i,j}$ into the reference frame coordinate space (current frame for which labels are generated) using vehicle ego-motion data.
III-B Local Object Motion Model
In our method, we classify object instances as either stationary or moving, as different principles of temporal consistency exploitation are applied to those classes. Note that all object instances actually appear to change position between frames – they change their location in relative terms, as the ego-vehicle itself is driving while data is captured.
In order to classify a vehicle as stationary or moving, we propose the following novel classification procedure.
For each instance, all locations $L_{m...n,j}$ are taken, and differences $\Delta_{i,j}$ between all adjacent locations are computed, where $k$ is the first frame, where the instance is present and $l$ is the last. From those differences, the standard deviation $\sigma$ is calculated as
$\sigma_{j}$ $=\frac{1}{\sqrt{2}}\sqrt{\frac{1}{l-k}}\sum_{i=k+1}^{l}(\mu_{j}-\Delta_{i,j})^{2}$ (4)
$\mu_{j}$ $=\frac{1}{l-k}\sum_{i=k+1}^{l}\Delta_{i,j}\qquad\Delta_{i,j}=L_{i,j}-L_{i-1,j}$ (5)
Furthermore, to decide if the difference between the $\left\lVert\mu_{j}\right\rVert_{2}$ and $\left\lVert\sigma_{j}\right\rVert_{2}$ is statistically significant or not, the ratio $z$ is calculated as
$z_{j}=\frac{\left\lVert\mu_{j}\right\rVert_{2}}{\left\lVert\sigma_{j}\right\rVert_{2}}$ (6)
To classify the instance as moving, the ratio $z$ is required to be higher than the threshold $T_{z}$ and simultaneously, the net distance must be higher than a given threshold $T_{m}$. This follows the idea, that if the car is stationary, the $\left\lVert\mu_{j}\right\rVert_{2}$ will be significantly lower than $\left\lVert\sigma_{j}\right\rVert_{2}$ as the movement of the vehicle is introduced by jitter from pseudo-LiDAR. On the other hand, if the vehicle is moving the $\left\lVert\mu_{j}\right\rVert_{2}$ will be significantly higher than $\left\lVert\sigma_{j}\right\rVert_{2}$.
As a result, for moving instances, we estimate the trajectory in adjacent frames. Given the known trajectory and the fact that cars move by physics constraints, estimating the yaw (the orientation) of the car is trivial. This simple fact simplifies the autolabelling significantly.
On the other hand, for the stationary instances, we can simply aggregate points $A_{j}$ by concatenating all $P_{i,j}$, as all $P_{i,j}$ are in the reference frame. This creates a much denser and more informative representation of an instance, and it can also recover from some mistakes, as shown in fig. 4.